# Dynamic Population Games: A Tractable Intersection of Mean-Field Games and Population Games

### v2

@@ IV-A Existence of SNE
Proof:
Consider the equivalent population game $F^{\mathcal{G}}(d,\pi)$ whose payoffs are given in Equations (9)–(10). Assumption 1 implies that $F^{\mathcal{G}}(d,\pi)$ is continuous in $(d,\pi)$: continuity of (9) is immediate; and continuity of (10) follows from noticing that the value function (5) can be equivalently written in vector from as $V_{\tau}(d,\pi)=\left(I-\alpha_{\tau}P_{\tau}(d,\pi)\right)^{-1}R_{\tau}(d,\pi)$, which is continuous in $(d,\pi)$ for all $\alpha_{\tau}\in[0,1)$. It follows that at least one NE $(d^{*},\pi^{*})$ is guaranteed to exist in $F^{\mathcal{G}}$ [20, Theorem 2.1.1], which coincides with a SNE of ${\mathcal{G}}$ by Theorem 1. ∎
IV-B Evolutionary Dynamics for SNE Computation
@@
where $H_{\tau}:{\mathbb{R}}^{|{\mathcal{A}}_{\tau}[x]|}\times\Delta({\mathcal{A}}_{\tau}[x])\rightarrow{\mathbb{R}}^{|{\mathcal{A}}_{\tau}[x]|}$ is one of the familiar mean dynamics in population games, such as the replicator, best response, Smith, or projection dynamic [20, Chapters 4–6], and $\eta_{\tau}$ is a rate parameter that controls the speed of policy dynamics (EV.2) with respect to state dynamics (EV.1). The dynamical system (EV.1)–(EV.2) is derived by applying evolutionary dynamics to the payoffs of the equivalent population game $F^{\mathcal{G}}(d,\pi)$ given in Equations (9)–(10). In particular, the projection dynamic is applied to (9) to yield (EV.1) [^1] In the projection dynamic, the social state is updated simply in the direction of the payoffs, which is projected onto the set of feasible directions keeping the social state in the simplex. Notice that the projection operation is redundant in this case.. This choice is interpretable since it yields Laplacian diffusion dynamics in (EV.1), however other mean dynamics could be used for the purpose of SNE computation. On the other hand, (EV.2) corresponds to applying mean dynamic $H_{\tau}$ to (10). For computation purposes, it is natural to choose a mean dynamic that satisfies the well-known condition of Nash stationarity [20, Section 5.2] for $H_{\tau}$, to guarantee that stationary points of $H_{\tau}$ coincide with NE of the underlying population game. Mean dynamics satisfying Nash stationarity include the best response, Smith, and projection dynamic (but not the replicator dynamic) [20, Table 5.1]. The following proposition is then immediate from Theorem 1.
Proposition 2
@@
A SNE can thus be computed by forward simulating dynamical system (EV.1)–(EV.2). If the system reaches a stationary point, a SNE is found. Notice that in comparison to existing mean field game approaches, the SNE computation is reduced to solving an ordinary differential equation (ODE) rather than an initial-terminal-value problem. Notice also that $\eta_{\tau}$, as well as the choice of the mean dynamic $H_{\tau}$ itself, are tuning parameters that can assist with the convergence to a SNE. For example, the Smith dynamic is better behaved than the best response dynamic in many examples of games [20]. Theorem 1 is also useful to provide formal convergence guarantees for (EV.1)–(EV.2), as discussed next.
IV-C Stability and Uniqueness of SNE
@@
Definition 2 is an immediate application of the definition of stable population games [20, Section 3.3] to the equivalent population game $F^{\mathcal{G}}$. Thus, the following proposition follows immediately from Theorem 1 and [20, Theorem 3.3.10, 7.2.1, 7.2.4, 7.2.7, 7.2.9].
Proposition 3
@@
If instead ${\mathcal{G}}$ is a strictly stable DPG, then it is guaranteed to have a unique SNE, which is globally asymptotically stable under the replicator, best response, Smith and projection dynamics.
V Applications
We present two application examples that demonstrate the versatility of our DPG formulation in tractably modelling complex real-world problems. The examples are adopted from recent application-specific works that have utilized the general results presented in this letter [7, 11, 13, 30, 31, 32].
V-A Fair Resource Allocation
@@
In all of the above works, the SNE is computed using the evolutionary dynamics model (EV.1)–(EV.1), with the perturbed best response dynamic [20] as the mean dynamic $H_{\tau}$. Figure 1 shows two representative examples of karma SNE adopted from [13], which correspond to a case where there are two agent types with different discount factors $\alpha_{1}=0.7$, $\alpha_{2}=0.99$. The top and bottom row of Figure 1 allow to assess the effect of a small progressive karma tax (collected from each agent and redistributed) on the SNE. The left column of Figure 1 shows the mean bid of the SNE policy $\pi^{*}$ ($y$-axis) per type and karma state ($x$-axis); the center column shows the stationary karma distribution $d^{*}$ per type; and the right column shows the stationary proportion of times each type gains access to the shared resource. With no tax, the future sighted agents bid slightly less (Figure 1 top-left) and end up holding more karma on average (Figure 1 top-center) than the short sighted agents; as a consequence, the future sighted agents are better off in terms of average access to the resource (Figure 1 top-right). The karma tax equalizes the field by redistributing karma to the short sighted agents. To the extent of our knowledge, such complex and high-dimensional SNE have not been computationally tractable before.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
V-B Epidemic Modelling and Control
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
VI Conclusion
We show that Stationary Nash Equilibria (SNE) in Dynamic Population Games (DPGs) can be reduced to standard Nash Equilibria (NE) in static population games. This result unlocks a rich set of existing tools for SNE analysis and computation, which makes the DPG formulation favorable in comparison to other mean-field game formulations. We demonstrate the versatility of the result in two complex application examples. An important future topic to investigate is the decentralized computation of the SNE by the agents.

### v3

@@ 4.1 Existence of SNE
Proof 4.4.
Consider the equivalent population game $F^{\mathcal{G}}(d,\pi)$ whose payoffs are given in Equations (9)–(10). Assumption 1 implies that $F^{\mathcal{G}}(d,\pi)$ is continuous in $(d,\pi)$: continuity of (9) is immediate; and continuity of (10) follows from noticing that the value function (5) can be equivalently written in vector from as $V_{\tau}(d,\pi)=\left(I-\alpha_{\tau}P_{\tau}(d,\pi)\right)^{-1}R_{\tau}(d,\pi)$, which is continuous in $(d,\pi)$ for all $\alpha_{\tau}\in[0,1)$. It follows that at least one NE $(d^{*},\pi^{*})$ is guaranteed to exist in $F^{\mathcal{G}}$ [25, Theorem 2.1.1], which coincides with a SNE of ${\mathcal{G}}$ by Theorem 1.
4.2 Evolutionary Dynamics for SNE Computation
@@
where $H_{\tau}$ is one of the familiar mean dynamics in population games, such as the replicator, best response, Smith, or projection dynamic, whose closed-form expressions are given in [25, Chapters 4–6]. The rate parameter $\eta_{\tau}$ controls the speed of policy dynamics (EV.2) with respect to state dynamics (EV.1). The dynamical system (EV.1)–(EV.2) is derived by applying evolutionary dynamics to the payoffs of the equivalent population game $F^{\mathcal{G}}(d,\pi)$ given in Equations (9)–(10). In particular, the projection dynamic is applied to (9) to yield (EV.1). This choice is interpretable since it yields Laplacian diffusion dynamics in (EV.1), however other mean dynamics could be used for the purpose of SNE computation. On the other hand, (EV.2) corresponds to applying mean dynamic $H_{\tau}$ to (10). For computation purposes, it is natural to choose a mean dynamic that satisfies the well-known condition of Nash stationarity [25, Section 5.2] for $H_{\tau}$, to guarantee that stationary points of $H_{\tau}$ coincide with NE of the underlying population game. Most existing mean dynamics satisfy Nash stationarity, including the best response, Smith, and projection dynamic (but not the replicator dynamic) [25, Table 5.1]. The following proposition is then immediate from Theorem 1.
Proposition 4.5.
@@
A SNE can thus be computed by forward simulating dynamical system (EV.1)–(EV.2). If the system reaches a stationary point, a SNE is found. Notice that discretizing (EV.1)–(EV.2) leads to many non-standard fixed point iterations of the SNE map (SNE.1)–(SNE.2) beyond the conventionally used Picard iteration [5, 22, 24]. For example, the best response dynamic leads to a Krasnoselskij iteration of (SNE.1)–(SNE.2), while other mean dynamics do not iterate the best response map (SNE.2) at all. The mean dynamic $H_{\tau}$, as well as the rate parameter $\eta_{\tau}$, are thus tuning parameters that can assist with the convergence to a SNE. For example, the Smith dynamic is better behaved than the best response dynamic in many examples of games [25]. Moreover, (EV.1)–(EV.2) enable adopting state of the art, higher-order learning techniques that have recently extended the classes of games for which these dynamics converge [26, 27, 28, 29, 30].
4.3 Stability and Uniqueness of SNE
@@
Definition 4.6 is an immediate application of the definition of stable population games [25, Section 3.3] to the equivalent population game $F^{\mathcal{G}}$. Thus, the following proposition follows immediately from Theorem 1 and [25, Theorem 3.3.10, 7.2.1, 7.2.4, 7.2.7, 7.2.9].
Proposition 4.7.
@@
If ${\mathcal{G}}$ is a strictly stable DPG, then it is guaranteed to have a unique SNE, which is globally asymptotically stable under the replicator, best response, Smith and projection dynamics.
5 Applications
We present two application examples that demonstrate the versatility of our DPG formulation in tractably modelling complex real-world problems. The modelling details of these examples are included in recent application-specific works that have utilized the general results of this letter [7, 11, 13, 32, 33, 34].
5.1 Fair Resource Allocation
@@
In all of the above works, the SNE is computed using the evolutionary dynamics (EV.1)–(EV.1), with $H_{\tau}$ chosen as the perturbed best response dynamic [25]. Figure 1 shows two representative examples of karma SNE adopted from [13], which correspond to a case where there are two agent types with different discount factors $\alpha_{1}=0.7$, $\alpha_{2}=0.99$. The top and bottom row of Figure 1 allow to assess the effect of a small progressive karma tax (collected from each agent and redistributed) on the SNE. The left column of Figure 1 shows the mean bid of the SNE policy $\pi^{*}$ ($y$-axis) per type and karma state ($x$-axis); the center column shows the stationary karma distribution $d^{*}$ per type; and the right column shows the stationary proportion of times each type gains access to the shared resource. With no tax, the future sighted agents bid slightly less (Figure 1 top-left) and end up holding more karma on average (Figure 1 top-center) than the short sighted agents; as a consequence, the future sighted agents are better off in terms of average access to the resource (Figure 1 top-right). The karma tax equalizes the field by redistributing karma to the short sighted agents. To the extent of our knowledge, such complex and high-dimensional SNE have not been computationally tractable before.
@@
5.2 Epidemic Modelling and Control
@@
6 Conclusion
We show that Stationary Nash Equilibria (SNE) in Dynamic Population Games (DPGs) can be reduced to standard Nash Equilibria (NE) in static population games. This result unlocks a rich set of existing tools for SNE analysis and computation, which makes the DPG formulation favorable in comparison to other mean-field game formulations. We demonstrate the versatility of the result in two complex application examples. Important future topics include investigating the relationship between condition (9) and previous conditions that guarantee convergence to a SNE, as well as the decentralized computation of the SNE by the agents.